If Benjamin is 4 years younger than Kevan and William is 4 years less than twice Benjamin's age and William is 22, how old are Kevan and Benjamin. What equation would I use?

Answer by sunny1992(35)

let
Benjamin = X year old
Kevan = X+4 year old
William = 2X-4 year old
given,
William = 22 year old
therefore,
2x-4=22
2x=22+4
2x=26
x=26/2
x=13
Benjamin = 13 year old
Kevan=13+4=17 year old